OpenStudy (anonymous):

Solve a=Πr^ 2 Would r=Πa^ 2 be the correct answer?

OpenStudy (tkhunny):

Seems unlikely at this point.

OpenStudy (anonymous):

There, I fixed it lol.

OpenStudy (tkhunny):

You do have the right idea, but it needs a little work. \(a = \pi r^{2} \rightarrow \dfrac{a}{\pi} = r^{2} \rightarrow \left(\dfrac{a}{\pi}\right)^{1/2} = r\)